A ship move from the port of Monrovia to the Port of Greenville and from there to the port of Bunchanan. The displacement vector is 10.2km due east, and the second is 5.9 km due norht. What is the resultant displacement vector?
1
Expert's answer
2020-09-08T09:12:09-0400

It can be found by Pythagorean theorem:


"R=\sqrt{d_E^2+d_N^2}=\\=\sqrt{10.2^2+5.9^2}=11.8\text{ km}."